what type of problem is this?: -3(2x+3)^2-6=-33
Answer by addingup(3677)

It's the type of problem you are supposed to solve.
Let's do it:
-3(2x+3)^2-6 = -33 add 6 to both sides:
-3(2x+3)^2 = -27 now divide both sides by -3, remember -/- = +
(2x+3)^2 = 9 take the square root of both sides
2x+3 = 3 OK, since we've been dealing with a quadratic you get 2 possible answers:
2x = -3-3 or 2x = 3-3
2x = -6 or 2x = 0
x = -3 or x = 0
Plug each into the original equation and see if you get -33.